OpenStudy (anonymous):

Take the derivative of log(base 2)x=(2^y)+(x^2)(y^2)

OpenStudy (anonymous):

no, dont take the derivative. be a non conformist. take the integral for gods sake!

OpenStudy (anonymous):

problem says take derivative

OpenStudy (anonymous):

haha ok you have two methods of doing this, one is easier than the other. the first is to just take the derivative, in which case you need to memorize a formula, and the other takes hardcore algebra. ill cover the first

OpenStudy (anonymous):

i have no idea O.o

OpenStudy (anonymous):

is the question to find dy/dx ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

needs both implicit and dealing with logs

OpenStudy (anonymous):

ok... do you know implicit differentiation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\huge log_2x=2^y+x^2y^2 \) \(\huge [log_2x]'=[2^y]'+[x^2y^2]' \) take the derivative of both sides... can you find all those derivatives?

OpenStudy (anonymous):

how do you take derivative of \[\log_{2}x \]

OpenStudy (anonymous):

1/xln2

OpenStudy (anonymous):

yes ^^^. you can get that by formula or just obtain it by change ofbase...

OpenStudy (anonymous):

how do you do the 2^y tho?

OpenStudy (anonymous):

i feel like i should know this -.- :/

OpenStudy (anonymous):

ok so for the 2 to the y do i also do ln

OpenStudy (anonymous):

\(\large [2^y]'=ln2 \cdot 2^y \cdot y' \) hehe... formula...

OpenStudy (anonymous):

... along with the chain rule of course...

OpenStudy (anonymous):

i feel embarrassed

OpenStudy (anonymous):

thats exactly what i had gotten, then ditched because i thought it was wrong

OpenStudy (anonymous):

that last one is using the product rule with the chain rule: \(\large [x^2y^2]'= \) ???

OpenStudy (anonymous):

so setting 2^y=w and then solving for dw= (you get 2^y * ln2 *y') is sufficiently how you would prove that formula for d 2^y ?

OpenStudy (anonymous):

yeah.. ^^^

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

so after u got all that... do some algebra isolate the y'....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yeah its simple implicit after that. id say the hardest part was recognizing that you can change the base in the 2^y.

OpenStudy (anonymous):

sorry... i have to make like a baby.... and head out..... cya... :)

OpenStudy (anonymous):

hwallday, you should have gotten something like ((1/xln2)-2xy^2)/((2^y)ln2+2yx^2)

OpenStudy (anonymous):

yup